# Training-Free Exponential Context Extension via Cascading KV Cache

### v2

@@ 1 Introduction
Large language models (LLMs) have become indispensable in a wide range of applications, from natural language processing to AI-driven content generation. However, their deployment is often hindered by the significant computational resources required, particularly during the quadratic attention operation in inference. Despite recent advancements like Flash Attention, which reduce memory overhead, the quadratic growth of latency and compute costs with input size remains a bottleneck, especially when processing long input sequences. This challenge is exacerbated in streaming applications, where high latency directly impacts user experience and operational costs.
@@
Existing methods, such as sliding window approaches (Beltagy et al., 2020; Jiang et al., 2023), attempt to manage long sequences but impose a static limit on the model’s ability to retain context due to the fixed window size. As a result, valuable tokens are naively discarded once they fall outside the window, leading to irreversible loss of context. While some recent methods have tried to stabilize this process by preserving a small number of initial tokens (Xiao et al., 2023) (Figure 3 top), they still operate within a rigid, static framework with a naive eviction policy that fails to account for the relative importance of tokens within the sequence. Additionally, the linear inference procedure from Xiao et al. (2023) processes a single token per step which makes it impractical to apply the method during the prompting stage when many tokens need to be processed together in parallel.
@@
Figure 2: Passkey retrieval accuracy up to 1M tokens given a cache size of 65K. Our Cascading cache maintains higher accuracy even after four doublings of the context length.
@@
Our work addresses these critical limitations by proposing a novel linear inference approach that extends the effective context length of sliding windows without increasing computational complexity or requiring additional training. We introduce a dynamic caching mechanism that organizes the key-value (KV) cache into cascading sub-caches, each designed to selectively retain tokens based on their historical importance. This approach allows for a more intelligent eviction policy, preserving tokens that are likely to play a crucial role in future predictions.
@@
To demonstrate the effectiveness of our method, we present two compelling examples: First, we visualize the attention matrix on the PG19 test set, showing how our cascading KV cache retains context far beyond the reach of a static sliding window, while maintaining the same total cache size (Figure 1). Second, we provide an example on passkey retrieval up to 1M tokens given a total cache size of 65K (Figure 2), illustrating that our method consistently outperforms static window attention in terms of retrieval accuracy, maintaining superior performance over prior work even after four doublings of the context size. These results underscore the potential of our approach to transform the efficiency and accuracy of LLMs in both research and real-world applications.
Our contributions are as follows:
@@
  • We provide a linear prefill strategy that avoids the restrictive quadratic prompt complexity of previous works. Our strategy reduces latency on 1M tokens by a factor of 6.8 compared to flash attention.
  • We provide an efficient implementation of our KV cache, which achieves a more than two order of magnitude speedup over Streaming LLM.
Figure 3: Comparison of Streaming LLM (Xiao et al., 2023) and Cascading Cache (Ours). Top: Streaming LLM stores fixed sink tokens (red) along with a sliding window of $N$ recent tokens. Bottom: Our method segments the cache into smaller cascading sub-caches, where each successive sub-cache conditionally accepts a fraction of tokens based on the magnitude of past attention scores. This simple technique allows for important tokens to remain in the cache for a longer time instead of being naively evicted too early.
@@ 2 Related Work
Our work builds on insights from Streaming LLM (Xiao et al., 2023), which identified a key phenomenon in transformers: as features progress through layers, attention scores increasingly concentrate on the initial tokens, referred to as ”sink tokens.” This phenomenon presents a challenge in traditional fixed sliding windows, as evicting sink tokens from the cache disrupts the attention score distribution, forcing the model to redistribute probability mass onto less ideal tokens. Xiao et al. (2023) demonstrated that this distribution shift leads to significant performance degradation unless the sink tokens are retained. We extend this insight by introducing an efficient, scalable method that increases the effective context window while maintaining a static cache size, improving throughput during the prompt stage, and enhancing performance on long-sequence tasks.
Subsequent approaches, such as SnapKV (Li et al., 2024) and H2O (Zhang et al., 2024), have introduced score-based KV cache eviction policies, but these focus primarily on quadratic prompts, with cache utilization limited to the decoding phase. These methods lack a linear prefill strategy, which results in the prefill attention operation remaining quadratic. Our method diverges from these previous approaches by offering a fast linear prefill strategy that is compatible with a cache that selectively retains important tokens through dynamic sparsity. This makes our approach not only more adaptive but also significantly more efficient, providing a practical path to extend transformer models’ context length with linear complexity and without incurring the high costs of retraining.
@@ 3 Method
Key-Value (KV) Caching. During LLM inference, a single token is generated at each time step. Combined with a causality condition such that token ${\bm{x}}_{i}$ cannot influence ${\bm{x}}_{j}$ iff $i>j$, it becomes more efficient to cache the $K^{(l)}$ and $V^{(l)}$ projected tokens in each layer $l$ rather than recomputing the full set of key, and value projections at each generation time step. Specifically, with $\cup$ representing a concatenation operation along the $S$ dimension, the calculation of the attention operation for the current token ${\bm{x}}_{j}$ during inference becomes,
${\bm{x}}_{j}^{(l+1)}=\sigma\left(\frac{1}{\sqrt{d}}\left({\bm{x}}_{j}^{(l)}{\bm{Q}}^{(l)}\right)\left({\bm{C}}^{(l)}_{K}\cup\left({\bm{x}}_{j}^{(l)}{\bm{K}}^{(l)}\right)\right)^{\top}\right)\left({\bm{C}}^{(l)}_{V}\cup\left({\bm{x}}_{j}^{(l)}{\bm{V}}^{(l)}\right)\right)\in{\mathbb{R}}^{d}$ (2)
After the concatenation operation, ${\bm{x}}_{j}^{(l)}{\bm{K}}^{(l)}$ and ${\bm{x}}_{j}^{(l)}{\bm{V}}^{(l)}$ are considered to be added to the ${\bm{C}}_{K}$ and ${\bm{C}}_{V}$ caches for the next iteration in the sequence.
@@
$\overline{{\bm{C}}}_{K}^{(t+1)}=\overline{{\bm{C}}}^{(t)}_{K}\cup{\bm{x}}_{j}{\bm{K}}=\left[\left\{{\bm{x}}_{i}{\bm{K}},{\bm{x}}_{i+1}{\bm{K}},\dots,{\bm{x}}_{j-1}{\bm{K}}\right\}\cup\left\{{\bm{x}}_{j}{\bm{K}}\right\}\right]\setminus\left\{{\bm{x}}_{i}{\bm{K}}\right\}.$ (3)
The only difference between standard sliding window attention and sink cache of Streaming LLM (Xiao et al., 2023) is that the sink cache perpetually retains the first $\alpha$ tokens in the sequence such that if a full cache starts from the first token ${\bm{x}}_{i}{\bm{K}}$, it is not token ${\bm{x}}_{i}{\bm{K}}$ which is evicted, but token ${\bm{x}}_{i+\alpha}{\bm{K}}$, with $\alpha\in{\mathbb{N}}$ being a fixed hyperparameter. Both sliding window attention and sink cache have the benefit of linear complexity, and a fixed overhead computation and memory cost for the window, as the size of the cache can only grow to a fixed, predetermined amount. However, this has the downside of constraining the information available in the cache, which may be needed for later predictions in the sequence. To illustrate, imagine streaming generation of an entire book with an LLM. If the cache can only contain tokens representing the average length of one chapter, then important information from previous chapters may be forgotten which could prove to be crucial for later steps of generation.
@@
3.1 Linear Prefill
A crucial limitation of Streaming LLM (Xiao et al., 2023), H2O (Zhang et al., 2024), and SnapKV (Li et al., 2024) lies in the handling of the prompt/prefill stage of the LLM. Each of the aforementioned works only considers that either 1) tokens are processed one-by-one during the prompt or 2) the prompt is processed with full quadratic attention and then proceeds to apply the relevant caching strategy during the generation phase. Being limited in this way, even a linear model like Streaming LLM exhibits slower latency than a quadratic prompt utilizing flash attention in the non-asymptotic regime (see Figure 6(b)). To bridge this gap, our method utilizes an attention kernel which processes fixed-sized chunks (strides) of the prompt in a single operation before adding the keys and values to our cascading cache. Specifically, for a sequence length of $S$, quadratic attention can be seen as having a stride size of $S$, Streaming LLM can be seen as having a stride size of $1$, and our prefill method can be seen as having a stride size $K\in[1,S]$ that is somewhere in between. Algorithm 1 contains pseudocode for our strided prefill process, and Figure 5 contains an illustration. Our strided prefill allows for a more than two order of magnitude decrease in latency compares to single token processing and reduces latency by a factor of $6.8$ over quadratic flash attention when processing 1M tokens. This significant improvement delivers the benefits of linear attention on realistic, non-asymptotic sequence lengths.
@@ 3.2 Cascading KV Cache
Positional Encoding. We use the same positional encoding strategy as Streaming LLM, which reapplies positional encodings for each token relatively and consecutively, always starting with the 0 index. With $\text{PE}_{\text{index}}:{\mathbb{N}}\mapsto{\mathbb{N}}$ signifying a positional encoding index mutation function, and $i,j\in{\mathbb{N}}$ representing the original token positions in the input sequence, it follows that $i<j\iff\text{PE}_{\text{index}}(i)<\text{PE}_{\text{index}}(j)$. For example, if our cascading cache holds the token indices from the original sequence $[0,1,3,5,7,8]$, they would in turn receive positional encodings $[0,1,2,3,4,5]$.
@@
Cache Token Span. Given previously outlined cascading cache, we have lengthened the span of the tokens which currently reside in the cache. The process outlined above effectively extends sliding window context length by allowing older tokens to remain as keys and values for a longer time, with gaps between tokens. Assuming that each sub-cache has the same capacity (i.e. $|{\bm{C}}|=|\bigcup_{i}^{N}{\bm{C}}_{i}|=N|{\bm{C}}_{1}|$), and each will accept tokens with a different frequency function defined as $\frac{1}{f(i)}$, then the approximate context length will be a summation over the cache sizes and the inverse of the frequency functions. For example, if each successive sub-cache accepts half of the tokens evicted from the previous cache, the total span of the cache $\tilde{S}$ becomes,
@@
We can then calculate the overall sparsity of the cache as $1-|C|/\tilde{S}$.
@@ 4 Experiments
Setup. We conduct experiments on streaming books (PG19) (Rae et al., 2019), Long Context Understanding (LongBench) (Bai et al., 2023b), book summarization (Booksum) (Kryściński et al., 2021), and 1M token passkey retrieval. We evaluate our method with pretrained transformers from the Llama3.1 (Dubey et al., 2024) and Qwen2 (Bai et al., 2023a) families of models. Please see Table 5 for model paths. In all our experiments, we keep the first 64 initial tokens as attention sinks. When considering the number of tokens in the cache, we always consider the sink tokens to be in addition to the cache size. Therefore a window size of $W$ has a total of $W+64$ tokens when accounting for the 64 sink tokens. We set the EMA parameter to $\gamma=0.9999$. We use the cascade token acceptance setting depicted in Figures 4 and 4, where each sub-cache accepts half of the tokens from the previous sub-cache. Unless otherwise indicated, our models use four cascading sub-caches. For experiments utilizing 8B model sizes, we use one NVIDIA A6000 (49GB) GPU, and for experiments utilizing 70B model sizes, we utilize 4 NVIDIA A100 GPUs. As Streaming LLM is a special case of our model, and 1 token per step is prohibitively slow, all results involving Streaming LLM results use our strided prefill strategy. Note that the strided prefill improves results over the original Streaming LLM with one token per step as shown in Figure 11.
@@
Figure 6: Latency. a) Our efficient cache implementation offers more than two orders of magnitude speedup over naive concatenation for cache add operations. b) Our strided prefill strategy is the only linear caching method which outperforms flash attention latency on realistic sequence lengths (1M tokens). We fit a second degree polynomial (dotted lines) to predict latencies after quadratic models run out of memory on a 49GB GPU. c) Strided prefill latency for 1M tokens by different stride sizes.
@@ 4.1 Latency
We compare the latency of our cascading cache to the implementation from Xiao et al. (2023) which uses tensor concatenation to add/evict tokens from the cache. Our implementation utilizes circular buffers and the Triton compiler (Tillet et al., 2019) to create an efficient CUDA kernel for the caching operation. To perform this experiment, we initialize a cache with 64 sink tokens and a total window size of $|{\bm{C}}|=16$K with 4 and 1 cascades, which are equivalent to our Cascading KV Cache and Streaming LLM, respectively. We also initialize the original Streaming LLM that uses concatenation. We then process a total of 16K tokens into the cache, and report the cumulative time spent on caching operations. Our method with one cascade (equivalent to Streaming LLM) takes just $0.01\%=1/10000$ of the total caching time of the original Streaming LLM, and our method with 4 cascades takes just $0.038\%=3.8/10000$ of the total caching time in Figure 6(a). In Figure 6(b), we show the overall attention latency, including caching, for a single attention layer processing 1M tokens. Our model uses a strided prefill of 4K with a total cache size of 16K. Our method is the only method which processes 1M tokens faster than flash attention, and takes only $14.8\%$ of the time of quadratic flash attention. In other words, flash attention is 6.8 times slower than ours for processing 1M tokens. We also study the effect of the size of our strided prefill on overall 1M token latency in Figure 6(b), finding that a stride of 4K delivers the lowest latency.

### v3

@@ 1 Introduction
Large language models (LLMs) have become indispensable in a wide range of applications, from natural language processing to AI-driven content generation. However, their deployment is often hindered by the significant computational resources required, particularly during the quadratic attention operation in inference. Despite recent advancements such as Flash Attention 2 (Dao, 2023), which reduce memory overhead, the quadratic growth of latency and compute costs with input size remains a burden, especially when processing long input sequences. This challenge is felt in streaming applications, where high latency directly impacts user experience and operational costs.
@@
Existing methods, such as sliding window approaches (Beltagy et al., 2020; Jiang et al., 2023), attempt to manage long sequences but impose a static limit on the model’s ability to retain context due to the fixed window size. As a result, valuable tokens are naively discarded once they fall outside the window, leading to irreversible loss of context. Although some recent methods have tried to stabilize this process by preserving a small number of initial tokens (Xiao et al., 2023) (Figure 3 top), they still operate within a static framework with a naive eviction policy that does not account for the relative importance of tokens within the sequence. Additionally, the linear inference procedure from Xiao et al. (2023) processes a single token per step, which is impractical to apply during the prompting stage when many tokens need to be processed together in parallel.
@@
Figure 2: Passkey accuracy up to 1M tokens given a cache size of 65K. Our Cascading cache maintains higher accuracy even after four doublings of the context length.
@@
Our work addresses these critical limitations by proposing a novel linear inference approach that extends the effective context length of sliding windows without increasing computational complexity or requiring additional training. We introduce a dynamic caching mechanism that organizes the key-value (KV) cache into cascading sub-caches, each designed to selectively retain tokens based on their importance. As opposed to a fixed sliding window that only evicts tokens at the end, our method offers multiple eviction routes before reaching the end, while intelligently preserving older tokens, which are likely to play a crucial role in future predictions.
@@
To demonstrate the effectiveness of our method, we present two compelling examples: First, we visualize the attention matrix on the PG19 test set, showing how our cascading KV cache retains context far beyond a static sliding window while maintaining the same total cache size (Figure 1). Second, we provide an example on passkey retrieval up to 1M tokens given a total cache size of 65K (Figure 2), illustrating that our method consistently outperforms static window attention in terms of retrieval accuracy, maintaining superior performance over prior work even after four doublings of the context size. These results underscore the potential of our approach to transform the efficiency and accuracy of long-context LLMs in both research and real-world applications. Our contributions are as follows:
@@
  • We provide a linear prefill strategy that avoids the restrictive quadratic prompt complexity of previous works. Our strategy reduces latency on 1M tokens by a factor of 6.8 compared to Flash Attention 2.
  • We provide an efficient implementation [^1] https://github.com/jeffwillette/cascading_kv_cache of our KV cache, which achieves a more than two order of magnitude speedup over Streaming LLM.
Figure 3: Comparison of Streaming LLM (Xiao et al., 2023) and Cascading Cache (Ours). Top: Streaming LLM stores fixed sink tokens (red) along with a sliding window of $N$ recent tokens. Bottom: Our method segments the cache into smaller cascading sub-caches, where each successive sub-cache conditionally accepts a fraction of tokens based on the magnitude of past attention scores. This simple technique allows for important tokens to remain in the cache for a longer time instead of being naively evicted too early. Conversely, superfluous tokens may be evicted before reaching the end of the cache, allowing for an intelligent eviction strategy.
@@ 2 Related Work
Our work builds on insights from Streaming LLM (Xiao et al., 2023), which identified a key phenomenon in transformers: as features progress through layers, attention scores often concentrate probability mass on the initial tokens, referred to as ”sink tokens.” This phenomenon presents a challenge in traditional fixed sliding windows, as evicting sink tokens from the cache disrupts the attention score distribution, forcing the model to redistribute probability mass onto other unexpected tokens. Xiao et al. (2023) demonstrated that this distribution shift leads to significant performance degradation unless the sink tokens are retained. We extend this insight by introducing an efficient, scalable method that increases the effective context window while maintaining a static cache size, improving throughput during the prompt stage, and enhancing performance on long-sequence tasks.
Subsequent approaches, such as SnapKV (Li et al., 2024) and H2O (Zhang et al., 2024b), have introduced score-based KV cache eviction policies, but these rely on quadratic prompts, with cache compression limited only to the decoding phase. These methods lack a linear prefill strategy, which results in the prefill attention operation remaining quadratic. Our method diverges from these previous approaches by offering a fast linear prefill strategy that is compatible with a cache that selectively retains important tokens through dynamic sparsity. This makes our approach not only more adaptive but also significantly more efficient, providing a practical path to extend transformer models’ context length with linear complexity and without incurring the high costs of retraining.
@@ 3 Method
Key-Value (KV) Caching. During LLM inference, a single token is generated at each time step. Combined with a causality condition such that token ${\bm{x}}_{i}$ cannot influence ${\bm{x}}_{j}$ iff $i>j$, it becomes more efficient to cache the $K^{(l)}$ and $V^{(l)}$ projected tokens in each layer $l$ rather than recomputing the full set of key, and value projections at each generation time step. Specifically, with $\cup$ representing a concatenation operation along the $S$ dimension, and each cache ${\bm{C}}\in\mathbb{R}^{S\times d}$, the calculation of the attention operation for the current token ${\bm{x}}_{j}$ during inference becomes,
${\bm{x}}_{j}^{(l+1)}=\sigma\left(\frac{1}{\sqrt{d}}\left({\bm{Q}}^{(l)}{\bm{x}}_{j}^{(l)}\right)^{\top}\left({\bm{C}}^{(l)}_{K}\cup\left({\bm{K}}^{(l)}{\bm{x}}_{j}^{(l)}\right)\right)^{\top}\right)\left({\bm{C}}^{(l)}_{V}\cup\left({\bm{V}}^{(l)}{\bm{x}}_{j}^{(l)}\right)\right)\in{\mathbb{R}}^{d}$ (2)
After the concatenation operation, ${\bm{x}}_{j}^{(l)}{\bm{K}}^{(l)}$ and ${\bm{x}}_{j}^{(l)}{\bm{V}}^{(l)}$ are considered to be added to the respective ${\bm{C}}_{K}$ and ${\bm{C}}_{V}$ caches for the next iteration in the sequence.
@@
$\overline{{\bm{C}}}_{K}^{(t+1)}=\overline{{\bm{C}}}^{(t)}_{K}\cup{\bm{K}}{\bm{x}}_{j}=\left[\left\{{\bm{K}}{\bm{x}}_{i},{\bm{K}}{\bm{x}}_{i+1},\dots,{\bm{K}}{\bm{x}}_{j-1}\right\}\cup\left\{{\bm{K}}{\bm{x}}_{j}\right\}\right]\setminus\left\{{\bm{K}}{\bm{x}}_{i}\right\}.$ (3)
The only difference between standard sliding window attention and sink cache of Streaming LLM (Xiao et al., 2023) is that the sink cache perpetually retains the first $\alpha$ tokens in the sequence such that if a full cache starts from the first token ${\bm{x}}_{1}{\bm{K}}$, it is not token ${\bm{x}}_{1}{\bm{K}}$ which is evicted, but token ${\bm{x}}_{1+\alpha}{\bm{K}}$, with $\alpha\in{\mathbb{N}}$ being a fixed hyperparameter. Both sliding window attention and sink cache have the benefit of linear complexity, and a fixed overhead computation and memory cost for the window, as the size of the cache can only grow to a fixed, predetermined amount. However, this has the downside of constraining the information available in the cache, which may be needed for later predictions in the sequence. To illustrate, imagine streaming generation of an entire book with an LLM. If the cache can only contain tokens representing the average length of one chapter, then important information from previous chapters may be forgotten which could prove to be crucial for later steps of generation.
@@
Linear Prefill. A crucial limitation of Streaming LLM (Xiao et al., 2023), H2O (Zhang et al., 2024b), and SnapKV (Li et al., 2024) lies in the handling of the prompt/prefill stage of the LLM. Each of the aforementioned works only considers that either 1) tokens are processed one-by-one during the prompt or 2) the prompt is processed with full quadratic attention and then proceeds to apply the relevant caching strategy during the generation phase. Being limited in this way, even a linear model like Streaming LLM exhibits slower latency than a quadratic prompt utilizing flash attention in the non-asymptotic regime (see Figure 6(b)). To bridge this gap, our method utilizes an attention kernel which processes fixed-sized chunks (strides) of the prompt in a single operation before adding the keys and values to our cascading cache. Specifically, for a sequence length of $S$, quadratic attention can be seen as having a stride size of $S$, Streaming LLM can be seen as having a stride size of $1$, and our prefill method can be seen as having a stride size $K\in[1,S]$ that is somewhere in between. Algorithm 1 contains pseudocode for our strided prefill process, and Figure 5 contains an illustration. Our strided prefill allows for a more than two order of magnitude decrease in latency compared to single token processing and reduces latency by a factor of $6.8$ over quadratic flash attention when processing 1M tokens. This significant improvement delivers the benefits of linear attention on realistic, non-asymptotic sequence lengths.
@@ 3.1 Cascading KV Cache
Positional Encoding. We use the same positional encoding strategy as Streaming LLM, which reapplies positional encodings for each token in the cache according to its index within the cache. Assuming that tokens with a higher index are more recently added to the cache, the positional encoding function $\text{PE}(\textbf{x}_{i},i)$ applies positional encoding index $i$ to token $\mathbf{x}_{i}$. For example, if our cache holds the token indices from the original sequence $[0,1,3,5,7,8]$, they would in turn receive positional encodings $[0,1,2,3,4,5]$ which corresponds to their index within the cache.
@@
Cache Token Span. Given previously outlined cascading cache, we have lengthened the span of the tokens which currently reside in the cache. The process outlined above effectively extends sliding window context length by allowing older tokens to remain as keys and values for a longer time, with gaps between tokens. Assuming that each sub-cache has the same capacity (i.e. $|{\bm{C}}|=|\bigcup_{i}^{N}{\bm{C}}_{i}|=N|{\bm{C}}_{1}|$), and each will accept tokens with a different frequency function defined as $\frac{1}{f(i)}$, then the approximate context length will be a summation over the cache sizes and the inverse of the frequency functions. For example, if each successive sub-cache accepts half of the tokens evicted from the previous cache, the total span of the cache $\tilde{S}$ becomes Equation 4. We can then calculate the overall sparsity of the cache as $1-|C|/S$ and sparsity of the cache window as $1-|C|/\tilde{S}$.
@@ 4 Experiments
Setup. We conduct experiments on streaming books (PG19) (Rae et al., 2019), Long Context Understanding (LongBench) (Bai et al., 2023b), book summarization (Booksum) (Kryściński et al., 2021), and 1M token passkey retrieval. We evaluate our method with pretrained transformers from the Llama3.1 (Dubey et al., 2024) and Qwen2 (Bai et al., 2023a) families of models. Please see Table 5 for model paths. In all our experiments, we keep the first 64 initial tokens as attention sinks. When considering the number of tokens in the cache, we always consider the sink tokens to be in addition to the cache size. Therefore a cache size of $|C|$ has a total of $|C|+64$ total tokens. We set the EMA parameter $\gamma=0.9999$. We use the cascade token acceptance policy depicted in Figures 4 and 4, where each sub-cache accepts half of the tokens from the previous sub-cache. Unless otherwise indicated, our models use four cascading sub-caches. As Streaming LLM is a special case of our model, and 1 token per step is prohibitively slow, all results involving Streaming LLM results use our strided prefill strategy. Note that the strided prefill improves results over the original Streaming LLM with one token per step as shown in Figure 11. See Appendix B for further details.
@@
Figure 6: Latency. a) Our efficient cache implementation offers more than two orders of magnitude speedup over naive concatenation for cache add operations. b) Our strided prefill strategy is the only linear caching method which outperforms flash attention 2 latency on realistic sequence lengths (1M tokens). We fit a second degree polynomial (dotted lines) to predict latencies after quadratic models run out of memory on a 49GB GPU. c) Strided prefill latency for 1M tokens by different stride sizes.
@@ 4.1 Latency
We compare the latency of our cascading cache to the implementation from Xiao et al. (2023) which uses tensor concatenation to add/evict tokens from the cache. Our implementation utilizes circular buffers and the Triton compiler (Tillet et al., 2019) to create an efficient CUDA kernel for the caching operation. To perform this experiment, we initialize a cache with 64 sink tokens and a total window size of $|{\bm{C}}|=16$K with 4 and 1 cascades, which are equivalent to our Cascading KV Cache and Streaming LLM, respectively. We also initialize the original Streaming LLM that uses concatenation. We then process a total of 16K tokens into the cache, and report the cumulative time spent on caching operations. Our method with one cascade (equivalent to Streaming LLM) takes just $0.01\%=1/10000$ of the total caching time of the original Streaming LLM, and our method with 4 cascades takes just $0.038\%=3.8/10000$ of the total caching time in Figure 6(a). In Figure 6(b), we show the overall attention latency, including caching, for a single attention layer processing 1M tokens. Our model uses a strided prefill of 4K with a total cache size of 16K. Our method is the only method which processes 1M tokens faster than flash attention 2, and takes only $14.8\%$ of the time of quadratic flash attention. In other words, flash attention is $6.8$ times slower than ours for processing 1M tokens. We also study the effect of the size of our strided prefill on overall 1M token latency in Figure 6(b), finding that a stride of 4K delivers the lowest latency.